joe has three children. The product of their ages is 36 and the sum of their ages is 21.What are the ages of Joe's children?
Answer by ankor@dixie-net.com(22740)

joe has three children.
The product of their ages is 36 and the sum of their ages is 21.
What are the ages of Joe's children?
:
Three kids: x, y, z
:
The product of their ages is 36
x*y*z = 36
:
the sum of their ages is 21.
x + y + z = 21
:
What are the ages of Joe's children?
:
Just looking at this I think the age of 1 child has to be 1
if x = 1
y*z = 36
and
y + z = 20
The only integers that work are 18 and 2
x = 1; y = 18; z = 2